A company mixes peanuts, cashews, and almonds to obtain 23 oz package worth 6.13 dollars. If peanuts, cashews, and almonds cost 11, 42, and 52 cents per oz respectively, and the weight of cashews in each package is equal to the weight of almonds, how much of each is in the package? ___oz of peanuts, ___oz of cashews, and ___oz of almonds.
Answer by TimothyLamb(4379)

x = oz of peanuts
y = oz of cashews
z = oz of almonds
---
x + y + z = 23
11x + 42y + 52z = 613
y = z
---
put the system of linear equations into standard form
---
x + y + z = 23
11x + 42y + 52z = 613
y - z = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 13
y = 5
z = 5
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---